OpenStudy (anonymous):

Given the linear equation y = -6x – 2, find the y-coordinates of the points (-1, ), (-2, ), and (1, ). Please show all of your work. Plot these points and graph the linear equation. Submit your graph through the dropbox.

OpenStudy (anonymous):

I guess no one knows how to set up this problem

OpenStudy (sswann222):

if the y corrdinates are (-1), (-2), and (1). Then you i believe you have to plug them in for x, then you are given a value for y for (-1) y= -6(-1) -2 = 6-2

OpenStudy (sswann222):

=4 for (-2) y = -6(-2) -2 = 12-2 = 10 for (1) y = 6(1)-2 = 6-2 = 4